two electrons are revolving around a nucleus at distances r and 4r. the ratio of their periods is

Grade:11

1 Answers

Vikas TU
7 years ago
Assuming the velocity of both the  electrons same.
T1 = 2pir/v
T2 = 2pi*4r/v
T1/T2 = r/4r = ¼
= > 1:4 would be the raatio of their Time periods.